OpenStudy (anonymous):

-17+√126x=√189x+9 x= this is the answer i got 29/(3√14-3√21) but my computer told me its wrong so im confused

OpenStudy (psymon):

-17 + 3sqrt(14)x = 3sqrt(21)x + 9x -17 = 3sqrt(21)x - 3sqrt(14)x + 9x -17 = [sqrt(21) - sqrt(14) + 3]3x \[x = \frac{ -17 }{ 3[3+\sqrt{21}-\sqrt{14}] }\]

OpenStudy (psymon):

I get something like this. Kinda see what i did?

OpenStudy (anonymous):

yes

OpenStudy (psymon):

Okay, awesome ^_^ Anything else you would do to that is unnecessary, so that seems like a well-simplified answer to me.

OpenStudy (anonymous):

well its not 9x its just 9

OpenStudy (psymon):

Oh xD It looks like 9x is all. Okay well then we would have instead: \[-26 = [\sqrt{21}-\sqrt{14}]3x\]= \[x=\frac{ -26 }{ 3(\sqrt{21}-\sqrt{14}) }\]

OpenStudy (anonymous):

thank you